OpenStudy (anonymous):

Which ordered pairs from the replacement set are solutions to the equation? y = -2x + 9 ; Replacement Set {(-1, 11), (0, 8), (1, 7), (2, 5)}

OpenStudy (anonymous):

@Easyaspi314

OpenStudy (anonymous):

Take each ordered pair and see if it satisfies the equation. I'll take the forst one, and you'll do the rest. The first ordered pair is (-1,11). Does -11 = -2(-1) + 9? -11 = 2 + 9? -11 = 11? No. So the first ordered pair does NOT work. Now you check the other ordered pairs.